# Higher derivative heterotic supergravity on a torus and supersymmetry

### v1

@@ 1 Introduction
While it is known that the $O(d,d+16;\mathbb{R})$ symmetry of the heterotic action persists to all orders in the $\alpha^{\prime}$ expansion, the actual structure of the higher-derivative couplings are not always apparent. A fruitful approach to constructing T-duality invariant actions is through double field theory Siegel:1993th ; Hull:2009mi ; Hohm:2010jy ; Hohm:2010pp and its higher derivative extensions. However, even in this case, the structure of the dimensionally reduced actions can be rather involved. Moreover, one generally needs to resort to field redefinitions in order to make the transformation properties manifest. At $\mathcal{O}(\alpha^{\prime})$, explicit $O(d,d;\mathbb{R})$ invariant dimensionally reduced bosonic and heterotic string effective actions were constructed in Eloy:2020dko . This construction involved a large set of non-trivial field redefinitions to put the action into a form that does not involve any additional derivatives on the fields beyond that expected from gauge covariance.
@@
The dimensional reduction makes use of the torsionful connections, $\Omega_{\pm}=\omega\pm\frac{1}{2}\mathcal{H}$, where $\omega$ is the conventional spin connection and $\mathcal{H}^{\alpha\beta}\equiv H_{\mu}{}^{\alpha\beta}dx^{\mu}$. This torsion connection arises naturally in the four-derivative heterotic action constructed in Bergshoeff:1988nn ; Bergshoeff:1989de , and is furthermore incorporated in higher derivative approaches to double field theory Marques:2015vua ; Baron:2017dvb and torus reductions David:2021jqn ; David:2022jcl . Using $\Omega_{\pm}$, we find that many terms in the reduction manifestly transform covariantly under $O(d)\times O(d)$. At the same time, there are a set of terms that do not, but a relatively straightforward $\mathcal{O}(\alpha^{\prime})$ field redefinition can be used to eliminate them.
@@ 4 Field redefinitions and removing explicit derivatives
$e^{-1}\mathcal{L}_{1}=e^{-2\phi}$ $\;\biggl{[}(R_{\alpha\beta\gamma\delta}(\omega_{+}))^{2}+\dfrac{1}{3}h_{\alpha\beta\gamma}\omega_{3L}(\omega_{+})_{\alpha\beta\gamma}-\dfrac{1}{2}R_{\alpha\beta\gamma\delta}(\omega_{+})F^{(-)\,a}_{\alpha\beta}F^{(-)\,a}_{\gamma\delta}$
@@
$e^{-1}\mathcal{L}_{2}=e^{-2\phi}$ $\;\biggl{[}-\dfrac{1}{8}F^{(+)\,a}_{\alpha\beta}F^{(+)\,a}_{\gamma\delta}F^{(+)\,b}_{\alpha\beta}F^{(+)\,b}_{\gamma\delta}+\dfrac{1}{4}F^{(+)\,a}_{\alpha\beta}F^{(+)\,b}_{\beta\gamma}F^{(+)\,a}_{\gamma\delta}F^{(+)\,b}_{\delta\alpha}$
@@
$e^{-1}\mathcal{L}_{3}=e^{-2\phi}$ $\;\biggl{[}\dfrac{1}{3}h_{\alpha\beta\gamma}\omega_{3}(-Q^{(++)})_{\alpha\beta\gamma}+\dfrac{1}{8}F^{(-)\,a}_{\alpha\beta}F^{(-)\,a}_{\gamma\delta}F^{(+)\,b}_{\alpha\beta}F^{(+)\,a}_{\gamma\delta}$
@@ 4.1 The Lagrangian with a torsion-free connection
$e^{-1}\mathcal{L}_{1}=e^{-2\phi}$ $\;\biggl{[}(R_{\alpha\beta\gamma\delta})^{2}-\dfrac{1}{2}R_{\alpha\beta\gamma\delta}F^{(-)\,a}_{\alpha\beta}F^{(-)\,a}_{\gamma\delta}-\dfrac{1}{8}(h_{\alpha\beta}^{2})^{2}$
@@
$e^{-1}\mathcal{L}_{2}=e^{-2\phi}$ $\;\biggl{[}-\dfrac{1}{32}F^{(+)\,a}_{\alpha\beta}F^{(+)\,a}_{\gamma\delta}F^{(+)\,b}_{\alpha\beta}F^{(+)\,b}_{\gamma\delta}-\dfrac{1}{8}F^{(+)\,a}_{\alpha\beta}F^{(+)\,a}_{\beta\gamma}F^{(+)\,b}_{\gamma\delta}F^{(+)\,b}_{\delta\alpha}$
@@
$e^{-1}\mathcal{L}_{3}=e^{-2\phi}$ $\;\biggl{[}-\dfrac{1}{4}F^{(-)\,a}_{\alpha\beta}F^{(-)\,a}_{\beta\gamma}F^{(+)\,b}_{\gamma\delta}F^{(+)\,b}_{\delta\alpha}-h^{2}_{\alpha\beta}P_{\alpha}^{(+-)\,ab}P_{\beta}^{(+-)\,ab}-\dfrac{1}{4}h^{2}_{\alpha\beta}F^{(+)\,a}_{\alpha\gamma}F^{(+)\,a}_{\beta\gamma}$
@@ 5 Discussion
Here we have only focused on heterotic supergravity. Had we included the ten-dimensional heterotic gauge fields, we would have had a $O(d,d+16)/O(d)_{-}\times O(d+16)_{+}$ coset and a corresponding enlarged flavor symmetry. Since we have written the reduced Lagrangian in a manifestly $O(d)_{+}\times O(d)_{-}$ notation, would be natural to simply extend all $O(d)_{+}$ indices to run from $1$ to $d+16$. While additional surprises may arise, we conjecture that this is sufficient to arrive at the $O(d)_{-}\times O(d+16)_{+}$ form of the reduced heterotic Lagrangian. This would be an example of symmetry leading the way to a more complete result with no additional work.

### v2

@@ 1 Introduction
While it is known that the $O(d,d+16;\mathbb{R})$ symmetry of the heterotic action persists to all orders in the $\alpha^{\prime}$ expansion, the actual structure of the higher-derivative couplings are not always apparent. A fruitful approach to constructing T-duality invariant actions is through double field theory Siegel:1993th ; Hull:2009mi ; Hohm:2010jy ; Hohm:2010pp and its higher derivative extensions Hohm:2013jaa ; Bedoya:2014pma ; Hohm:2014xsa ; Coimbra:2014qaa ; Lee:2015kba ; Marques:2015vua ; Baron:2017dvb ; Lescano:2021guc . However, even in this case, the structure of the dimensionally reduced actions can be rather involved. Moreover, one generally needs to resort to field redefinitions in order to make the transformation properties manifest. At $\mathcal{O}(\alpha^{\prime})$, explicit $O(d,d;\mathbb{R})$ invariant dimensionally reduced bosonic and heterotic string effective actions were constructed in Eloy:2020dko ; Elgood:2020xwu ; Ortin:2020xdm . This construction involved a large set of non-trivial field redefinitions to put the action into a form that does not involve any additional derivatives on the fields beyond that expected from gauge covariance.
@@
The dimensional reduction makes use of the torsionful connections, $\Omega_{\pm}=\omega\pm\frac{1}{2}\mathcal{H}$, where $\omega$ is the conventional spin connection and $\mathcal{H}^{\alpha\beta}\equiv H_{\mu}{}^{\alpha\beta}dx^{\mu}$. This torsion connection arises naturally in the four-derivative heterotic action constructed in Bergshoeff:1988nn ; Bergshoeff:1989de , and is furthermore incorporated in higher derivative approaches to double field theory Marques:2015vua ; Baron:2017dvb and torus reductions Elgood:2020xwu ; Ortin:2020xdm ; David:2021jqn ; David:2022jcl . Using $\Omega_{\pm}$, we find that many terms in the reduction manifestly transform covariantly under $O(d)\times O(d)$. At the same time, there are a set of terms that do not, but a relatively straightforward $\mathcal{O}(\alpha^{\prime})$ field redefinition can be used to eliminate them.
@@ 4 Field redefinitions and removing explicit derivatives
$e^{-1}\mathcal{L}_{1}=e^{-2\varphi}$ $\;\biggl{[}(R_{\alpha\beta\gamma\delta}(\omega_{+}))^{2}+\dfrac{1}{3}h_{\alpha\beta\gamma}\omega_{3L}(\omega_{+})_{\alpha\beta\gamma}-\dfrac{1}{2}R_{\alpha\beta\gamma\delta}(\omega_{+})F^{(-)\,a}_{\alpha\beta}F^{(-)\,a}_{\gamma\delta}$
@@
$e^{-1}\mathcal{L}_{2}=e^{-2\varphi}$ $\;\biggl{[}-\dfrac{1}{8}F^{(+)\,a}_{\alpha\beta}F^{(+)\,a}_{\gamma\delta}F^{(+)\,b}_{\alpha\beta}F^{(+)\,b}_{\gamma\delta}+\dfrac{1}{4}F^{(+)\,a}_{\alpha\beta}F^{(+)\,b}_{\beta\gamma}F^{(+)\,a}_{\gamma\delta}F^{(+)\,b}_{\delta\alpha}$
@@
$e^{-1}\mathcal{L}_{3}=e^{-2\varphi}$ $\;\biggl{[}\dfrac{1}{3}h_{\alpha\beta\gamma}\omega_{3}(-Q^{(++)})_{\alpha\beta\gamma}+\dfrac{1}{8}F^{(-)\,a}_{\alpha\beta}F^{(-)\,a}_{\gamma\delta}F^{(+)\,b}_{\alpha\beta}F^{(+)\,a}_{\gamma\delta}$
@@ 4.1 The Lagrangian with a torsion-free connection
$e^{-1}\mathcal{L}_{1}=e^{-2\varphi}$ $\;\biggl{[}(R_{\alpha\beta\gamma\delta})^{2}-\dfrac{1}{2}R_{\alpha\beta\gamma\delta}F^{(-)\,a}_{\alpha\beta}F^{(-)\,a}_{\gamma\delta}-\dfrac{1}{8}(h_{\alpha\beta}^{2})^{2}$
@@
$e^{-1}\mathcal{L}_{2}=e^{-2\varphi}$ $\;\biggl{[}-\dfrac{1}{32}F^{(+)\,a}_{\alpha\beta}F^{(+)\,a}_{\gamma\delta}F^{(+)\,b}_{\alpha\beta}F^{(+)\,b}_{\gamma\delta}-\dfrac{1}{8}F^{(+)\,a}_{\alpha\beta}F^{(+)\,a}_{\beta\gamma}F^{(+)\,b}_{\gamma\delta}F^{(+)\,b}_{\delta\alpha}$
@@
$e^{-1}\mathcal{L}_{3}=e^{-2\varphi}$ $\;\biggl{[}-\dfrac{1}{4}F^{(-)\,a}_{\alpha\beta}F^{(-)\,a}_{\beta\gamma}F^{(+)\,b}_{\gamma\delta}F^{(+)\,b}_{\delta\alpha}-h^{2}_{\alpha\beta}P_{\alpha}^{(+-)\,ab}P_{\beta}^{(+-)\,ab}-\dfrac{1}{4}h^{2}_{\alpha\beta}F^{(+)\,a}_{\alpha\gamma}F^{(+)\,a}_{\beta\gamma}$
@@ 5 Discussion
Here we have only focused on heterotic supergravity. Had we included the ten-dimensional heterotic gauge fields, we would have had a $O(d,d+16)/O(d)_{-}\times O(d+16)_{+}$ coset and a corresponding enlarged flavor symmetry. Since we have written the reduced Lagrangian in a manifestly $O(d)_{-}\times O(d)_{+}$ notation, would be natural to simply extend all $O(d)_{+}$ indices to run from $1$ to $d+16$. While additional surprises may arise, we conjecture that this is sufficient to arrive at the $O(d)_{-}\times O(d+16)_{+}$ form of the reduced heterotic Lagrangian. This would be an example of symmetry leading the way to a more complete result with no additional work.